Hayhayz (hayhayz):

Andrea can clean a house 4 times as fast as Denise. When they work together, Andrea and Denise can clean a large house in 8 hours. How many hours would it take Denise to clean the house by herself? Answer choices : 20, 15, 10 , 5

OpenStudy (escamer):

@malcolmmcswain

OpenStudy (escamer):

u know he knows this haleigh ;) XD

OpenStudy (escamer):

@dan815

Hayhayz (hayhayz):

@dan815

OpenStudy (anonymous):

Let x = Denise's time if do by herself 4x = Andrea's time if do by herself the time when they work together can be written as \[(4x * x)/(4x + x) = 8\] \[4x^2/5x = 8\] \[4x/5 = 8\] \[4x = 5 * 8\] \[4x = 40\] \[x = 10\] so, Denise's time id work by herself is 10 hours

Hayhayz (hayhayz):

thank youu

OpenStudy (dan815):